OpenStudy (anonymous):

vector{3,4,5}, how to count the vector length? And the sequences of the number represent x,z,y axis or wat?

OpenStudy (anonymous):

yeah d given vector means : 3i + 4j + 5k

OpenStudy (anonymous):

wat does i ,j and k mean?

OpenStudy (anonymous):

the magnitude of x y and z axis...

OpenStudy (anonymous):

so in my case x=3,z=4 and y=5? and how do u count the vector length?

OpenStudy (anonymous):

nup...x=3 y=4 n z=5

OpenStudy (mathmate):

Vector length |(3,4,5)| = sqrt(3^2+4^2+5^2)=sqrt(9+16+25)=sqrt(50)=5sqrt(2)

OpenStudy (anonymous):

ty guys ^^

OpenStudy (anonymous):

one more thing, if vector{3,4,5} represent a rectangle in x,y,x, wat about vector{3,4,5,6}? Wat does the 6 mean? and any brief idea about how the shape is?

OpenStudy (mathmate):

Unfortunately there is no visual image of a 4 dimensional object. Usually the fourth dimension is time.

OpenStudy (mathmate):

Some mathematicians describe the object in four dimensions as a hypercube, but we cannot have a hypercube physically.

OpenStudy (anonymous):

vector can have infinity of variable in it?